# Graph Neural Networks for O-RAN Mobility Management: A Link Prediction Approach

### v2

@@ Abstract
Mobility performance has been a key focus in cellular networks up to 5G. To enhance Handover (HO) performance, 3GPP introduced Conditional Handover and Layer 1/Layer 2 Triggered Mobility (LTM) mechanisms in 5G. While these reactive HO strategies address the trade-off between HO failures and ping-pong effects, they often result in inefficient radio resource utilization due to additional HO preparations. To overcome these challenges, this article proposes a proactive HO framework for mobility management in O-RAN, leveraging user-cell link predictions to identify the optimal target cell for HO. We explore various categories of Graph Neural Networks (GNNs) for link prediction and analyze the complexity of applying them to the mobility management domain. Two GNN models are compared using a real-world dataset, with results demonstrating their ability to capture the dynamic and graph-structured nature of cellular networks. Finally, we present key insights from our study and outline future steps to enable the integration of GNN-based link prediction for mobility management in O-RAN networks.
Index Terms:
Graph Neural Networks, O-RAN, Handover, Mobility Management, Link Prediction.
@@ I Introduction
Mobility management is a fundamental feature of cellular networks, ensuring seamless service continuity during user mobility by minimizing call drops, Radio Link Failures (RLFs) and unnecessary handovers (HOs), known as ping-ponging [1]. This management can be carried out through different HO mechanisms, which are responsible for decision-making and messages handling in the RAN. Traditionally, mobility management follows a reactive HO mechanism, where HO decisions rely on predefined network parameters. However, determining optimal mobility parameters remains complex due to the trade-off between HO failures and ping-pong effects. Additionally, the dynamic communication environment often limits reactive techniques. These challenges are further exacerbated in 5G networks, where densification, higher frequencies, and stringent URLLC requirements demand more sophisticated mobility management solutions [2].
@@
The advent of Open Radio Access Network (O-RAN) has revolutionized the RAN by leveraging Artificial Intelligence (AI) or Machine Learning (ML) to enhance network performance and flexibility. These AI/ML models, deployed at the edge and centralized cloud, enable intelligent decision-making and automation in managing modern cellular networks. Mobility management is a key O-RAN use case, and recent studies have explored proactive HO mechanisms based on user trajectory [3], user measurements [4], and RLF prediction [5]. ML models such as sequence-to-sequence (Seq2Seq) and Long Short-Term Memory (LSTM) show promise in capturing long-term dependencies between user and cell metrics. However, these models process data sequentially, failing to leverage the graph-structured nature of communication networks, limiting generalization and reducing predictive performance on unseen networks and mobility patterns.
@@
To address the challenges of modeling dynamic network topologies, Graph Neural Networks (GNNs) have become increasingly relevant in both academic and industrial context [6, 7]. GNNs operate on graph-structured data by iteratively aggregating information from neighboring nodes, making them suitable for tasks such as node or edge regression and link prediction. Prior work has applied GNNs to regression problems in mobility management. such as predicting signal quality or load-related metrics [8]. Although these outputs can support HO optimizations, the models stop at intermediate predictions, with the estimation of HO targets left as an unexplored work.
@@
In this article, we study the capabilities of GNN link prediction for mobility management and how to overcome the technical challenges when adapting these ML models as proactive HO mechanism. This work is intended as an initial exploration, aimed at demonstrating the feasibility of using GNN-based link prediction to anticipate HOs in O-RAN systems. We propose next-cell prediction, a proactive HO mechanism that leverages a GNN-based link prediction model to forecast the upcoming cell connection for each user. We provide an in-depth analysis and model comparison of two categories of GNN link prediction models and analyze the complexity of applying them to the mobility management problem. Sec. II reviews the O-RAN framework, current HO mechanisms and the role of GNNs, while also summarizing recent advancements in proactive HO mechanisms. Sec. III outlines the problem formulation, proposes an O-RAN framework for next-cell prediction, and analysis in-depth the challenges of applying GNNs for mobility management. Sec. IV compares the performance of prediction accuracy of subgraph-based and autoencoder-based GNN link prediction models, trained on real-world datasets. Finally, Sec. V discusses the key findings and insights on the applicability of GNN-based link prediction for mobility management, concluding the article with practical implications and future directions.
@@ II-A O-RAN
The O-RAN Alliance aims to create an intelligent, open, multi-vendor, interoperable, and autonomous radio access network. A core innovation of O-RAN is the introduction of Radio Intelligent Controllers (RICs). These include the non-Real-Time (non-RT) RIC for long-term optimization and planning, hosted within the Service Management and Orchestration (SMO). Complementing this is the near-Real-Time (near-RT) RIC for dynamic and real-time network configuration. O-RAN further fragments the 3GPP architecture, which already splits the Central Unit (CU) and the Distributed Unit (DU), by additionally dividing the CU into Control Plane (CU-CP) and User Plane (CU-UP).
@@ II-B Handover mechanisms in O-RAN
This work addresses inter-gNB HO in O-RAN, utilizing the Xn interface for transitions between different gNBs, as shown in Fig. 1. The baseline HO process involves the User Equipment (UE) measuring neighboring cell metrics, such as Signal-to-Noise Ratio (SNR), and reporting these to its source gNB (1). If the source signal degrades and neighbor’s improves, the source gNB initiates the HO procedure. This involves a sequence of steps: an XnAP Handover Request is sent to the target gNB (2), the target gNB replies with an acknowledgment (3), the source gNB then instructs the UE to reconfigure its connection (4), a temporary user-plane tunnel is set up between source and target CU-UPs for data transfer, and finally the UE confirms the reconfiguration to the target gNB (6).
O-RAN facilitates the integration of advanced AI/ML-based solutions in the RICs via rApps and xApps. The distinction lies in their response time restrictions: rApps (non-RT RIC) operate with response times greater than 1 second, while xApps (near-RT RIC) require response times between 10 ms and 1 second. In the standard [9], O-RAN introduces the use of these AI/ML solutions to implement proactive HO mechanisms for mobility management. The non-RT RIC trains the AI/ML models using collected metrics via the O1 interface and then the deployed rApps sends A1 policies to the near-RT RIC. The xApp deployed in the near-RT RIC subsequently uses these A1 policies, alongside E2 measurements, to apply real-time configuration updates, facilitating proactive and adaptive HO decisions.
@@ II-D Existing solutions for mobility management
Recent AI/ML approaches for mobility management primarily use Seq2Seq models for optimizing intra-frequency HO based on UE trajectories [3], and LSTM models for incorporating mobility information into next-cell prediction [4]. While these methods show promising results, they rely on sequential data processing using Recurrent Neural Networks (RNN) or Reinforcement Learning (RL) approaches. These architectures are limited in their ability to capture the structural relationships present in cellular networks and often struggle to generalize across evolving topologies [10].
Cellular networks naturally form graphs of interdependent nodes with features, making GNNs suitable for solving network problems. They have been applied to radio resource management [2], 5G signal inference [11] and traffic prediction [6]. In the case of link prediction, models must account for factors like interference and load, which introduce temporal variability. Recent studies combine GNNs with RL, as in [8], where connection management is performed as a graph optimization task. While expressive, RL-based methods are often data-intensive and computationally expensive, limiting real-time applicability. Industry efforts such as Amazon’s GNN-based mobility solution [7] highlight the practical interest in this domain, though technical details remain unpublished. Furthermore, [5] combines GNNs and time-series transformers to predict RLFs, but focuses on link reliability than mobility management.
@@ III-B Next-Cell prediction rApp - A Walkthrough
The proposed O-RAN mobility management solution is shown on Fig. 2 and operates by integrating the next-cell prediction rApp into the Non-RT RIC. The workflow starts with UE collecting and transmitting measurement reports (1). Subsequently, both the serving and neighboring gNBs transmit RAN data and configurations, including L1/L2/L3 metrics, to the SMO Operations, Administration, and Maintenance (OAM) functions via the O1 interface (2). These OAM functions have a baseline measurement configuration to select the data of interest, and allows its access to the Non-RT RIC (3). A pre-trained GNN-based link prediction model, hosted as an rApp evaluates the collected RAN data to forecast the upcoming cell connections. Based on these predictions, the rApp generates A1 Policy update messages and transmits them to the Connection Management xApp via the A1 interface (4). This xApp, based on the the A1 policies and the E2 measurement metrics obtained from the E2 nodes (5.a and 5.b), generates RIC Control messages (6) to initiate radio resource re-allocation and HO request (7 and 8), mirroring the standard HO process from Fig. 1. Crucially, the only HO latency introduced by this solution is attributed to O1 data collection, model inference (evaluated in Sec. II-D), and policy generation, all of which are low-overhead operations. This design choice is fundamental to maintaining efficiency and adhering to the stringent timing requirements inherent in O-RAN deployments.
@@ IV Model evaluation
This section evaluates two GNN models: VGAE based and SEAL. VGAE captures global embeddings, while SEAL focuses on localized subgraphs, providing a complementary comparison for mobility management. The following subsections detail the dataset, model creation, training methodology, and experimental setup.
IV-A Dataset description
The study utilizes an open-source dataset from Huawei Technologies [13], derived from a real-world operational network. It initially comprises approximately 35 million edges and 169,000 nodes (100,000 user nodes and 69,000 cell nodes). For privacy, all node and edge features are anonymized, represented as random properties normalized between 0 and 1, which limits the data analysis.
The data is structured across three main files:
  • node_cell: Contains 69,000 cell IDs, each with 1,000 associated feature columns
  • node_user: Includes 100,000 user IDs, each with 300 feature columns
  • edge_user_live_cell: Originally lists about 35 million edges, each specifying an edge ID, user ID, cell ID, and 37 edge-specific features. To manage potential redundancy and due to a lack of information on multiple edges, only one edge per user-cell pair (the one with the lowest ID) was used, reducing the total edges to approximately 7 million.
This dataset presents each user connected to 69 cells and each cell to 1,000 users, creating 1,000 isolated sub-networks without inter-connectivity. This structure yields an extremely sparse network density of 0.00049. Due to computational complexity, initial model development and feasibility studies were conducted on a subset of 985 cells and 10,000 users, with plans for gradual scaling. This specific dataset was selected for its scale, structure, and operational origin, providing a strong foundation for an initial GNN-based link prediction model development and feasibility study in O-RAN. Further validation across different datasets and network scenarios will be necessary to fully assess generalization performance.
@@ IV-B Model creation
The link prediction task is formally defined over an attributed graph $G=(V,E,X),$ where $V$ denotes the set of nodes, $E$ the set of edges, and $X$ represents node and edge features. The objective is to predict the probability of a link existing between pairs of nodes $(u,v)$, corresponding to the likelihood that a user will handover to a given target cell.
Model adaptations were necessary due to the graph’s specific nature and the solution’s objectives. As the dataset comprises two node types connected through a single hop, the model design employs a single layer for data transmission between neighbors. Incorporating additional layers would result in redundant information transmission and lead to over-smoothing and learning inefficiencies. To address each challenges of GNN models for mobility management, the following adaptations are implemented:
@@ IV-C Model training
A systematic training process was applied to optimize model performance. The dataset was divided into 80/10/10 training, validation, and test sets using a custom split function to ensure consistent, structure-preserving partitions. This split generates positive (existing) and negative (non-existing) edges for each set, which are necessary for contrastive supervision during training. This custom approach prevents potential issues associated with automatic split generation functions (e.g., PyTorch RandomLinkSplit), such as including existing edges as negative, and duplicated negative edges between different sets, which negatively impact model performance [14] [15].
@@ IV-D Performance comparison
Fig. 3 shows the mean metrics values obtained from 10 executions of each model. Both models, SEAL (blue) and VGAE (light orange), demonstrate comparable predictive performance over the 985 cells and 10,000 UEs graph. VGAE improves SEAL’s AP and Recall by 6.21% and 11.23%. Conversely, SEAL outperformed VGAE in AUC (4.5%), Precision (12.3%), F1-score (10.09%), Accuracy (0.61%), and MCC (2.81%). While VGAE effectively identifies edges within a broader context, it yields more False Positives (FP), thus excelling in FP-independent metrics like Recall. However, SEAL surpasses VGAE in other metrics due to VGAE’s higher FP count. SEAL’s use of localized subgraph information also enhanced its precision in predictions with smaller datasets.
@@
The two darker orange bars in Fig. 3 examine VGAE scalability by increasing the graph size to 10,000 cells and 20,000 cells, both with the 100,000 UEs. SEAL could not run on these larger datasets due to the significant computational cost associated with subgraph creation. VGAE performance improved overall with the increase of the dataset size, from VGAE-985 to VGAE-10k. Recall showed highest growth (4.91%), followed by AUC (3%), Accuracy (2.2%) and F1-score (2.18%). This reflects an increase in correctly predicted edges, due to improved generalization from training the model on a broader variety of subgraph structures and interactions. As a result, the model learns to assign more accurate attention to nodes during training. However, VGAE-20k model performance deteriorates. This occurred because the dataset maintained the same number of UEs, but duplicated cells, leading to a drastic increase in node density, which likely affects the attention mechanism and increases FPs (key limitation of the model).
@@
Fig. 4 shows the models training times. VGAE outperforms SEAL in execution speed, achieving a 98.84% reduction in training time with the 985 cells dataset due to its direct graph processing. This computational burden prevented SEAL from executing on the two large datasets. VGAE also demonstrates faster convergence during training, achieving satisfactory results within less than 50 epochs. This rapid convergence, coupled with its simpler architecture, enables efficient training, even on larger datasets. Furthermore, VGAE’s training time increases linearly with the graph size as it performs embeddings for each node of the network.
@@
Fig.5 illustrates VGAEs reduction in SEAL’s inference time with the 985 cells dataset, also attributed to its direct execution without subgraphs. VGAEs inference time, similar to its training time, also grows linearly with dataset size. Crucially, all model inference times are below or close to 1 second, demonstrating their high efficiency and suitability for O-RAN rApp deployments, which are characterized by operational response times greater than 1 second.
@@
Figure 3: Performance metrics obtained with VGAE and SEAL, along with different dataset sizes.
@@
Figure 4: Training time of VGAE and SEAL, along with different dataset sizes.
@@
Figure 5: Inference time of VGAE and SEAL, along with different dataset sizes..

### v3

@@ Abstract
Mobility performance has been a key focus in cellular networks up to 5G. To enhance Handover (HO) performance, 3GPP introduced Conditional Handover and Layer 1/Layer 2 Triggered Mobility (LTM) mechanisms in 5G. While these reactive HO strategies address the trade-off between HO failures and ping-pong effects, they often result in inefficient radio resource utilization due to additional HO preparations. To overcome these challenges, this article proposes a proactive HO framework for mobility management in O-RAN, leveraging user-cell link predictions to identify the optimal target cell for HO. We explore various categories of Graph Neural Networks (GNNs) for link prediction and analyze the complexity of applying them to the mobility management domain. Two GNN models are compared using an anonymized real-world dataset and a synthetic dataset obtained with Keysight’s EXata emulation tool. The results show the models’ ability to capture the dynamic and graph-structured nature of cellular networks. Finally, we present key considerations for real-world deployment that outline future steps to enable the integration of GNN-based link prediction for mobility management in O-RAN networks.
@@ I Introduction
Mobility management is a fundamental feature of cellular networks, ensuring seamless service continuity during user mobility by minimizing call drops, Radio Link Failures (RLFs) and unnecessary handovers (HOs), known as ping-ponging [1]. This management can be carried out through different HO mechanisms, which are responsible for decision-making and message handling in the Radio Access Network (RAN). Traditionally, mobility management follows a reactive HO mechanism, where HO decisions rely on predefined network parameters. However, determining optimal mobility parameters remains complex due to the trade-off between HO failures and ping-pong effects. Additionally, the dynamic communication environment often limits reactive techniques. These challenges are further exacerbated in 5G networks, where densification, higher frequencies, and stringent URLLC requirements demand more sophisticated mobility management solutions [2].
@@
The advent of Open Radio Access Network (O-RAN) has revolutionized the RAN by leveraging Artificial Intelligence (AI) or Machine Learning (ML) to enhance network performance and flexibility. These AI/ML models, deployed at the edge and centralized cloud, enable intelligent decision-making and automation in managing modern cellular networks. Mobility management is a key O-RAN use case, and recent studies have explored proactive HO mechanisms based on user trajectory [3] and User Equipment (UE) measurements [4]. ML models such as sequence-to-sequence (Seq2Seq) and Long Short-Term Memory (LSTM) show promise in capturing long-term dependencies between UE and cell metrics. However, these models process data sequentially, failing to leverage the graph-structured nature of communication networks, limiting generalization and reducing predictive performance on unseen networks and mobility patterns.
@@
To address the challenges of modeling dynamic network topologies, Graph Neural Networks (GNNs) have become increasingly relevant in both academic and industrial contexts [5, 6]. GNNs operate on graph-structured data by iteratively aggregating information from neighboring nodes, making them suitable for tasks such as node or edge regression and link prediction. Prior work has applied GNNs to regression problems in mobility management, such as predicting signal quality or load-related metrics [7]. Although these outputs can support HO optimizations, the models stop at intermediate predictions, with the estimation of HO targets remaining unexplored.
@@
In this article, we study the capabilities of GNN link prediction for mobility management and how to overcome the technical challenges when adapting these ML models as proactive HO mechanism. This work is intended as an initial exploration, aimed at demonstrating the feasibility of using GNN-based link prediction to anticipate HOs in O-RAN systems. We propose next-cell prediction, a proactive HO mechanism that leverages a GNN-based link prediction model to forecast the upcoming cell connection for each user. We provide an in-depth analysis and model comparison of two categories of GNN link prediction models and analyze the complexity of applying them to the mobility management problem. Sec. II reviews the O-RAN framework, current HO mechanisms and the role of GNNs, while also summarizing recent advancements in proactive HO mechanisms. Sec. III outlines the problem formulation, proposes an O-RAN framework for next-cell prediction, and analyzes in-depth the challenges of applying GNNs for mobility management. Sec. IV compares the performance of prediction accuracy of subgraph-based and autoencoder-based GNN link prediction models, trained on real-world and synthetic datasets. Sec. V summarizes the main considerations for the real-world deployment of GNN-based link prediction for mobility management. Finally, Sec. VI concludes the article with key findings and future directions.
@@ II-A O-RAN
O-RAN is an industry effort to open and disaggregate the RAN, enabling multi-vendor interoperability, software driven control, and AI/ML assisted automation. Its architecture centers on two RAN Intelligent Controllers (RICs), the Non-Real-Time (Non-RT) RIC hosted in the Service Management and Orchestration (SMO) system, and the Near-Real-Time (Near-RT) RIC close to the RAN. rApps run in the Non-RT to train models and issue policies over the A1 interface, while xApps run in the Near-RT RIC to apply those policies together with measurements received over the E2 interface. Network configuration and performance data flow between the SMO and RAN through the O1 interface. In parallel, functional disaggregation separates the Central Unit (CU) and the Distributed Unit (DU) and further splits the CU into control plane (CU-CP) and user plane (CU-UP) parts, which simplifies scaling, upgrades, and vendor mixing within a gNB deployment.
@@ II-B Handover mechanisms in O-RAN
This work addresses inter-gNB HO in O-RAN, utilizing the Xn interface for transitions between different gNBs, as shown in Fig. 1. The baseline HO process involves the UE measuring neighboring cell metrics, such as Signal-to-Noise Ratio (SNR), and reporting these to its source gNB (1). If the source signal degrades and neighbor’s improves, the source gNB initiates the HO procedure. This involves a sequence of steps: an XnAP Handover Request is sent to the target gNB (2), the target gNB replies with an acknowledgment (3), the source gNB then instructs the UE to reconfigure its connection (4), a temporary user-plane tunnel is set up between source and target CU-UPs for data transfer, and finally the UE confirms the reconfiguration to the target gNB (6).
O-RAN facilitates the integration of advanced AI/ML-based solutions in the RICs via rApps and xApps. The distinction lies in their response time restrictions: rApps operate with response times greater than 1 second, while xApps require response times between 10 ms and 1 second. In the standard  [8], O-RAN introduces the use of these AI/ML solutions to implement proactive HO mechanisms for mobility management. The Non-RT RIC trains the AI/ML models using collected metrics via the O1 interface and then the deployed rApp sends A1 policies to the Rear-RT RIC. The xApp deployed in the near-RT RIC subsequently uses these A1 policies, alongside E2 measurements, to apply real-time configuration updates, facilitating proactive and adaptive HO decisions.
@@ II-D Existing solutions for mobility management
Recent AI/ML approaches for mobility management [9] include Seq2Seq models for intra-frequency HO optimizations using UE trajectories [3], LSTM models for next-cell prediction with mobility context [4] and Deep Reinforcement Learning (DRL) methods that learn policies from interactions [10, 9]. These families of models capture temporal patterns and can adapt to changing traffic and load. In this article, we study a graph-based formulation that operates on the network structure, with the goal of assessing whether link prediction can support proactive decisions in an O-RAN. We propose this as complementary to DRL and recurrent models, as each family targets different aspects of the problem.
Cellular networks naturally form graphs of interdependent nodes with features, making GNNs suitable for solving network problems. They have been applied to radio resource management [2], 5G signal inference [11] and traffic prediction [5]. In the case of link prediction, models must account for factors like interference and load, which introduce temporal variability. Recent studies combine GNNs with RL, as in [7], where connection management is performed as a graph optimization task. Industry efforts such as Amazon’s GNN-based mobility solution [6] highlight the practical interest in this domain, though technical details remain unpublished. Furthermore, [12] combines GNNs and time-series transformers to predict RLFs, but focuses on link reliability than mobility management.
@@ III-B Next-Cell prediction rApp - A Walkthrough
The proposed O-RAN mobility management solution is shown on Fig. 2 and operates by integrating the next-cell prediction rApp into the Non-RT RIC. The workflow starts with UE collecting and transmitting measurement reports (1). Subsequently, both the serving and neighboring gNBs transmit RAN data and configurations, including L1/L2/L3 metrics, to the SMO Operations, Administration, and Maintenance (OAM) functions via the O1 interface (2). These OAM functions have a baseline measurement configuration to select the data of interest, and allows its access to the Non-RT RIC (3). A pre-trained (with prior knowledge of the network) GNN-based link prediction model, hosted as an rApp evaluates the collected RAN data to forecast the upcoming cell connections. Based on these predictions, the rApp generates A1 Policy update messages and transmits them to the Connection Management xApp via the A1 interface (4). This xApp, based on the the A1 policies and the E2 measurement metrics obtained from the E2 nodes (5.a and 5.b), generates RIC Control messages (6) to initiate radio resource re-allocation and HO request (7 and 8), mirroring the standard HO process from Fig. 1. Crucially, the only HO latency introduced by this solution is attributed to O1 data collection, model inference (evaluated in Sec. IV.D), and policy generation, all of which are low-overhead operations. This design choice is fundamental to maintaining efficiency and adhering to the stringent timing requirements inherent in O-RAN deployments.
@@ IV Model evaluation
This section evaluates two GNN models: VGAE based and SEAL. VGAE captures global embeddings, while SEAL focuses on localized subgraphs, providing a complementary comparison for mobility management. The following subsections detail the , model creation, training methodology, and experimental setup.
IV-A Used datasets
To validate the model performance we used two datasets: i) an open-source Real-World dataset (RW dataset) from Huawei Technologies [^1] https://snap.stanford.edu/data/telecom-graph.html, and ii) a synthetic EMulated dataset (EM dataset) generated using Keysight 5G network emulator EXata [^2] https://www.keysight.com/fr/en/product/SN100VGLA/exata-network-modeling-5g.html. RW was selected for its scale, structure, and operational origin, providing a strong foundation for an initial GNN-based link prediction model development. The small-scale EXata dataset (EM), was generated for validation of the models generalization capabilities across diverse network scenario with real radio metrics. The datasets main characteristics are:
  • RW is composed of 35 million edges and 169,000 nodes (100,000 UEs and 69,000 cells), which allows the analysis of the models’ scalability. All nodes and edge have features associated: UEs have 1,000, cells have 300, and edges have 37. However, for privacy reasons, all node and edge features are anonymized, represented as random properties normalized between 0 and 1, which limits the data analysis.
  • EM is composed of 449,152 edges and 101 nodes (70 UEs and 31 cells). The emulator allows the collection of L1/L2 UE radio metrics. Performing feature engineering on the collected data and excluding the highly correlated features, as UE-cell edges we used: RSRP, RSRQ, Transport Blocks, Available Resources Blocks, Packet Size, MCS, Signal Power Uplink, and Signal Power Downlink.
From the raw data, we create the same graph data structure for both datasets. Due to computational complexity of processing RW and the scale differences, initial development is conducted on a subset of 985 cells and 10,000 users, with further for gradual scaling. Also, RW limits its interpretability due to a lack of information on multiple edges between the same pair of UE-cell. To manage potential redundancy, only one edge per user-cell pair (the one with the lowest ID) was used, reducing the total edges to approximately 7 million. Same procedure is then used in EM, reducing the edges to 489. This way, RW has each UE connected to 69 cells and each cell to 1,000 UEs, creating 1,000 isolated sub-networks without inter-connectivity. EM has each UE connected with different number of cells, ranging from two to six. These structures yields sparse network densities of 0.00049 and 0.09683, respectively.
@@ IV-B Model creation
The link prediction task is formally defined over an attributed graph $G=(V,E,X),$ where $V$ denotes the set of nodes, $E$ the set of edges, and $X$ represents node and edge features. The objective is to predict the probability of a link existing between pairs of nodes $(u,v)$, corresponding to the likelihood that a user will HO to a given target cell.
Model adaptations were necessary due to the graph’s specific nature and the solution’s objectives. As the dataset comprises two node types connected through a single hop, the model design employs a single layer for data transmission between neighbors. Incorporating additional layers would result in redundant information transmission and lead to over-smoothing and learning inefficiencies.
To address each challenge of GNN models for mobility management, the following adaptations are implemented:
@@ IV-C Model training
A systematic training process was applied to optimize model performance. The datasets was divided into 80/10/10 training, validation, and test sets using a custom split function to ensure consistent, structure-preserving partitions. This split generates positive (existing) and negative (non-existing) edges for each set, which are necessary for contrastive supervision during training. This custom approach prevents potential issues associated with automatic split generation functions (e.g., PyTorch RandomLinkSplit), such as including existing edges as negative, and duplicated negative edges between different sets, which negatively impact model performance [15].
@@ IV-D Performance comparison
Fig. 3 shows the mean metrics values obtained from 10 executions of each model-dataset combination. The blue bars correspond to the results with the EM dataset with SEAL (SEAL-EM) and VGAE (VGAE-EM). The orange bars represent SEAL and VGAE applied to the smallest partition of RW, with 985 cells and 10,000 UEs (SEAL-985-RW and VGAE-985-RW), and the VGAE results when scaling the dataset to 10,000 cells and 100,000 UEs (VGAE-10k-RW) and to 20,000 cells and 1,000,000 UEs (VGAE-20k-RW).
Using the EM dataset, VGAE-EM outperforms SEAL-EM in AUC (0.81%), AP(18.63%), Precision (6.08%), F1-score (0.1%) and Accuracy (1.36%), whereas SEAL-EM outperforms VGAE-EM in Recall (7.53%). Meanwhile, using the 985-RW dataset, VGAE-985-RW achieves higher AP (6.21%) and Recall (11.23%) than SEAL-985-RW, but SEAL-985-RW outperforms VGAE-985-RW in AUC (4.5%), Precision (12.3%), F1-score (10.09%) and Accuracy (0.61%). The differing outcomes observed in the models’ performance, depending on the used dataset, arise from their inherent differences in density and size. SEAL shows lower performance with the EM, as the low dataset size produces small subgraphs that limit its learning capacity. In contrast, VGAE enhances its performance with EM, as it is almost 200 times denser than RW, which particular benefits to autoencoder-based models capturing edges within a broader context. This is the reason for VGAE worsening with RW, generating more False Positives (FP) due to the sparsity of the graph. However, with the larger dataset SEAL achieves stronger performance in FP-sensitive metrics.
@@
Fig. 3 also illustrates VGAE scalability with increasing graph size. SEAL could not run on these larger datasets due to the significant computational cost associated with subgraph creation. VGAE performance improved overall when scaling from VGAE-985-RW to VGAE-10k-RW, with Recall showing the highest growth (4.91%), followed by AUC (3%), Accuracy (2.2%) and F1-score (2.18%). This reflects an increase in correctly predicted edges, due to improved generalization from training the model on a broader variety of subgraph structures and interactions. As a result, the model learns to assign more accurate attention to nodes during training. However, VGAE-20k-RW model performance deteriorates. This occurred because the dataset maintained the same number of UEs, but duplicated cells, leading to a drastic increase in node density, which likely affects the attention mechanism and increases FPs (key limitation of the model).
@@
Fig. 4 shows the models training times. VGAE outperforms SEAL in execution speed, achieving a 98.84% reduction in training time for 985-RW and a 38.5% for EM due to its direct graph processing. This computational burden prevented SEAL from executing on the two large datasets. VGAE also demonstrates faster convergence during training, achieving satisfactory results within less than 50 epochs in both datasets. This rapid convergence, coupled with its simpler architecture, enables efficient training, even on larger datasets. Furthermore, VGAE’s training time increases linearly with the graph size as it performs embeddings for each node of the network.
@@
Fig.5 illustrates VGAEs reduction in SEAL’s inference time for 985-RW and for EM, also attributed to its direct execution without subgraphs. VGAEs inference time, similar to its training time, also grows linearly with dataset size. Crucially, all model inference times are below or close to 1 second, demonstrating their high efficiency and suitability for O-RAN rApp deployments, which are characterized by operational response times greater than 1 second.
@@
Figure 3: Performance metrics obtained with VGAE and SEAL, along with different dataset types and sizes.
@@
Figure 4: Training time of VGAE and SEAL, along with different dataset types and sizes.
@@
Figure 5: Inference time of VGAE and SEAL, along with different dataset types and sizes.